OpenStudy (ali2x2):

Please solve the following inequality, showing all work and then list 3 numbers that are in the solution set. x/6 ≥ 3 In your own words: - Tell how to solve inequalities. Be specific and complete. - Tell how to graph inequalities. Be specific and complete. Please solve the following inequality, showing all work and then list 3 numbers that are in the solution set. x/6 + 4 ≥ 2

OpenStudy (ali2x2):

Hello?

OpenStudy (astrophysics):

Hi, well what have you tried?

OpenStudy (astrophysics):

\[\frac{ x }{ 6 } \ge 3\] this should be an easy one, how would you solve for x?

OpenStudy (ali2x2):

Im not smart when it comes to this :/ @Astrophysics

OpenStudy (astrophysics):

That's ok, just try your best I don't judge :)

OpenStudy (astrophysics):

How would you go about isolating x in your question, it's just like normal algebra with a funny looking sign, that has a few extra rules, such as when you divide or multiply by negative, the sign flips :P

OpenStudy (ali2x2):

well i thought placeing 6 on the so it would be 6/6 = 1 but the 3 wins and that wouldnt be right

OpenStudy (ali2x2):

:P

OpenStudy (astrophysics):

Hey do you know how to use the draw tool?

OpenStudy (ali2x2):

MHm

OpenStudy (astrophysics):

 just show what you mean on here

OpenStudy (astrophysics):

You can click on the drawing on top right, there should be a pencil

OpenStudy (ali2x2):